A random sample of 100 students at southern
university was collected. The mean age was found to be 24.2 years, with a standard deviation of 3.4 years
A) what percent of students are over 27.6 years old
b)what percent of the students are between 20.8 and 24.2 years old
c) If a student is selected at random from southern univesity, what is the probability that he or she will be over 20.8 years old


I could use some help

Answer by stanbon(48569)

A random sample of 100 students at southern
university was collected. The mean age was found to be 24.2 years, with a standard deviation of 3.4 years
A) what percent of students are over 27.6 years old
z(27.6) = (27.6-24.2)/[3.4] = 1
P(x > 27.6) = P(z > 1) = 0.15866
----------------------------------------------------------------
b)what percent of the students are between 20.8 and 24.2 years old
z(20.8) = (20.8-24.2)/3.4 = -1
z(24.2) = 1
So, P(20.8 < x < 24.2) = P(-1 < z < 1) = 0.68269
------------------------------------------------------------------
c) If a student is selected at random from southern univesity, what is the probability that he or she will be over 20.8 years old
p(x > 20.8) = P(z > -1) = 0.84134
=========================================
Cheers,
Stan H.